# Sequential Change Detection with Differential Privacy

### v1

@@ IV Proposed $\delta$-DP-CUSUM Procedure with Unbounded Log-Likelihood Ratio
$\mathbf{P}_{\tilde{T},X_{k},X^{\prime}_{k}}(\tilde{T}=t\mid X_{(1:t)\setminus k})\leq e^{\epsilon}\mathbf{P}_{\tilde{T},X_{k},X^{\prime}_{k}}(\tilde{T}=t\mid X^{\prime}_{(1:t)\setminus k})+\delta,$ (16)
@@
We would like to emphasize that this relaxed $(\epsilon,\delta)$-differential privacy guarantee in (16) is not the same as the classical definition of $(\epsilon,\delta)$-DP in literature [9], which considers the probability taken over the randomness in the random algorithm while the database is fixed. Here in this slightly modified version (16), we use a small probability $\delta$ to capture the failure case when the LLR at differing entries, $|\ell(X_{k})-\ell(X^{\prime}_{k})|$, is greater than $A_{\delta}$; we will still achieve $\epsilon$-DP conditional on this failure event not occurring, and $\delta$ accounts for such failures. To avoid confusion, we named our proposed procedure as $\delta$-DP-CUSUM procedure instead of $(\epsilon,\delta)$-DP-CUSUM procedure.
@@
For the general case with unbounded LLR, for any $\epsilon>0$, $\delta\in(0,1)$ and the corresponding $A_{\delta}$ defined in (15), we have
@@
Then, for any $\epsilon>0$, $\delta\in(0,1)$ and the corresponding $A_{\delta}$ defined in (15), we have the worst-case average detection delay satisfies
@@
$b=b_{\gamma}=\frac{\log\gamma+O(\log\log\gamma)}{h(\epsilon,A_{\delta})}=\frac{\log\gamma}{h(\epsilon,A_{\delta})}(1+o(1)),$ (19)
to guarantee the ARL constraint $\mathbf{E}_{\infty}[\tilde{T}]\geq\gamma$ is met. Under this choice of threshold, the resulting WADD of the $\delta$-DP-CUSUM procedure is upper bounded by
@@
Our proposed procedure will achieve the first-order asymptotical optimality as the classical CUSUM procedure if $\epsilon\geq 2A_{\delta}.$ On the other hand, if $\epsilon<2A_{\delta}$, the upper bound becomes $\frac{\log\gamma}{\epsilon/2A_{\delta}\cdot I_{0}}(1+o(1))$, which is greater than the optimal detection delay $\frac{\log\gamma}{I_{0}}(1+o(1))$ of classical CUSUM. Therefore, a smaller value of $h$ implies greater degradation in detection performance due to the privacy requirement. We further provide a Gaussian example below to illustrate how the privacy parameters $\epsilon$ and $\delta$ affect the detection efficiency of our proposed procedure by visualizing the value of $h(\epsilon,A_{\delta}).$ We emphasize that $\delta$ is pre-specified by the user and it yields the trade-off between privacy stringency and detection delay. In practice, one can choose the largest $\delta$ acceptable under one’s privacy requirements to make $h$ closer to 1, and thus reduce the detection delay.
@@
We provide an example under Gaussian distributions to demonstrate the theoretical effect of privacy parameters $\epsilon,\delta$ on detection delay. Assume one-dimensional Gaussian mean shift from the pre-change distribution $N(0,1)$ to the post-change distribution $N(\mu,1)$ with $\mu>0$. Therefore $f_{0}(x)=\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}x^{2}}$ is the pdf of pre-change distribution, and $f_{1}(x)=\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}(x-\mu)^{2}}$ is the pdf of post-change distribution. The LLR equals $\ell(X)=\log f_{1}(X)/f_{0}(X)=\mu X-\frac{1}{2}\mu^{2}$. Taking $A_{\delta}=2|\mu|\cdot z_{\delta/4}+\mu^{2}$ ensures $\max_{i=0,1}\,\mathbf{P}_{X\sim f_{i}}\left(2|\ell(X)|\geq A_{\delta}\right)\leq\delta/2$. The quantity $h(\epsilon,A_{\delta})=\min\{\frac{\epsilon}{2A_{\delta}},1\}$ is the effective privacy scaling factor that directly influences the asymptotic performance of the $\delta$-DP-CUSUM procedure. In particular, the best-case scenario corresponds to $h(\epsilon,A_{\delta})=1$, under which the $\delta$-DP-CUSUM procedure is guaranteed to be first-order asymptotically optimal. In contrast, smaller values of $h$ indicate a greater degradation in detection efficiency. Fig. 1 visualizes $h(\epsilon,A_{\delta})$ over a wide range of $\epsilon$ and $\delta$ values for different magnitude of the mean shift $\mu\in\{0.1,0.25,0.5\}$. The plots show that $h(\epsilon,A_{\delta})$ generally decreases as $\mu$ increases, since $A_{\delta}$ increases with $\mu$ and thus leading to a stronger impact of privacy on detection performance. The red dashed line in Fig. 1 serves as the boundary for optimal detection given by Theorem 2: any privacy parameters satisfying $\epsilon\geq 2A_{\delta}$ (equivalently $h(\epsilon,A_{\delta})=1$) guarantee that the $\delta$-DP-CUSUM procedure achieves first‑order asymptotic optimality.
@@ V-A DP-CUSUM with Bounded LLR: Laplace Distribution
We consider a mean shift in Laplace distribution family $\text{Laplace}(\mu,1),$ which has a pdf $f(x|\mu)=\exp\left(-|x-\mu|\right)/2.$ Note we use a different notation here because we have used the shorthand notation $\text{Lap}(b)$ before for denoting the zero‑mean Laplace noise added in the detection procedure. We let the pre-change distribution be $\text{Laplace}(0,1)$ and the post-change distribution be $\text{Laplace}(\mu,1)$, with a nonzero post-change mean $\mu\neq 0$. Note the LLR between $\text{Laplace}(\mu,1)$ and $\text{Laplace}(0,1)$ is $\ell(x)=\left(|x|-|x-\mu|\right),$ which implies the sensitivity of the LLR is $\Delta=\sup_{x,x^{\prime}\in\mathbb{R}}|\ell(x)-\ell(x^{\prime})|=2|\mu|$.
@@
In Fig. 2(a), it can be seen that the detection delays for $\epsilon=0.8$ and $1$ are close to that of the exact CUSUM. This is consistent with our theoretical results on WADD in Theorem 3. Specifically, for $\mu=0.2$, we have $\Delta=2\mu=0.4$. By Eq. (14), the DP-CUSUM procedure is guaranteed to be asymptotically optimal when $\epsilon\geq 2\Delta=0.8$, which aligns with Fig. 2(a). As $\epsilon$ decreases, the delay increases, especially at larger ARL values, reflecting the tradeoff between privacy and detection. A similar pattern is observed in Fig. 2(b) for $\mu=0.5$, where $\Delta=1$. It can be seen that the delay for $\epsilon=2\Delta=2$ matches the exact CUSUM, while smaller $\epsilon$ leads to worse performance.
@@
(a) $\text{Laplace}(0,1)$ $\to$ $\text{Laplace}(0.2,1)$ (b) $\text{Laplace}(0,1)$ $\to$ $\text{Laplace}(0.5,1)$
@@
We further compare the performance of our DP-CUSUM procedure with the baseline method, online private change-point detector (OnlinePCPD) [25], as shown in Fig. 3. OnlinePCPD uses a fixed-size sliding window over the data stream. Within each window, it performs offline private change-point detection based on CUSUM statistics. Since the detection statistic is recomputed in each window, it is non-recursive and incurs a higher computational cost. Notably, in OnlinePCPD, Laplace noise $\text{Lap}(4\Delta/\epsilon)$ is added to the detection statistic and $\text{Lap}(8\Delta/\epsilon)$ to the threshold. In contrast, our method only adds $\text{Lap}(2\Delta/\epsilon)$ to both the statistic and the threshold, while still achieving the same level of $\epsilon$-DP, as guaranteed by Theorem 1. We adopt the same window size of 700 as used in [25]. Fig. 3 shows that the detection delay of OnlinePCPD is substantially larger than that of our DP-CUSUM procedure across all privacy regimes (even when $\epsilon\leq 2\Delta$), particularly when the ARL is large, under both mean shift scenarios $\mu=0.2$ and $\mu=0.5$ in the Laplace distribution setting.
@@
(a) $\text{Laplace}(0,1)$ $\to$ $\text{Laplace}(0.2,1)$ (b) $\text{Laplace}(0,1)$ $\to$ $\text{Laplace}(0.5,1)$
@@ V-B $\delta$-DP-CUSUM with Unbounded LLR: Normal Distribution
For $\mu=0.1$, we have $A_{\delta}=0.402$. By Eq. (20), the $\delta$-DP-CUSUM procedure is guaranteed to be asymptotically optimal when $\epsilon\geq 2A_{\delta}=0.804$, which aligns with the results in Fig. 4(a) where the detection delays for $\epsilon=1$ and $1.5$ are close to that of the exact CUSUM. A similar pattern is observed for $\mu=0.5$, where $A_{\delta}=2.21$. In Fig. 4(b), the delay for $\epsilon=4$ is close to that of the exact CUSUM, while smaller $\epsilon$ leads to worse performance. It is also worth noting that performance degradation is less pronounced for smaller $\mu$ under the same $\epsilon$, likely because the detection problem is already difficult for smaller $\mu$ values—leading to larger delays even for the exact CUSUM procedure.
@@ VI Conclusion
We introduce a theoretically grounded DP-CUSUM procedure for sequential change detection with differential privacy constraints. We derive the lower bound to the average run length and the upper bound to the worst-case average detection delay of DP-CUSUM. The theoretical results also yield the first-order asymptotic optimality of DP-CUSUM under weak privacy requirements. Experimental results on synthetic data demonstrate the good performance of our DP-CUSUM procedure compared with the baseline method and confirm our theoretical findings. Our theoretical analysis provides analytical techniques for future extensions and the design of DP-CUSUM variants across diverse practical applications. Future work includes adapting this framework to more complex settings, such as unknown post-change distributions, multi-sensor detection with controlled sensing, network-structured data, and related applications.
@@ Appendix A Useful Lemmas
$|S_{t}-S^{\prime}_{t}|=|\max\left(0,S_{t-1}\right)+\ell(X_{t})-\max\left(0,S^{\prime}_{t-1}\right)-\ell(X_{t})|\leq|S_{t-1}-S^{\prime}_{t-1}|\leq\Delta_{t-1}.$
@@
Lemma 4 ([31, 18]).
@@
The proof can be found in [31, Lemma 3] and we omit here. ∎
@@
$\mathbf{E}_{\infty}[e^{\lambda S_{t}}]$ $=\int_{0}^{\infty}\mathbf{P}_{\infty}(e^{\lambda S_{t}}\geq y)dy=\int_{0}^{\infty}\mathbf{P}_{\infty}(S_{t}\geq\frac{\log y}{\lambda})dy$
$\overset{(i)}{=}\int_{-\infty}^{\infty}\mathbf{P}_{\infty}(S_{t}\geq u)\lambda e^{\lambda u}du=\int_{-\infty}^{0}\mathbf{P}_{\infty}(S_{t}\geq u)\lambda e^{\lambda u}du+\int_{0}^{\infty}\mathbf{P}_{\infty}(S_{t}\geq u)\lambda e^{\lambda u}du$
@@
where $(i)$ by change of variables $y=e^{\lambda u}$, and $(ii)$ by $\mathbf{P}_{\infty}[S_{t}\geq u]\leq 1$ and $\mathbf{P}_{\infty}(S_{t}\geq u)\leq e^{-u}$.

### v2

@@ IV Proposed $\delta$-DP-CUSUM Procedure with Unbounded Log-Likelihood Ratio
$\mathbf{P}_{\tilde{T},X_{k},X^{\prime}_{k}}(\tilde{T}=t\mid X_{(1:t)\setminus k})\\ \leq e^{\epsilon}\mathbf{P}_{\tilde{T},X_{k},X^{\prime}_{k}}(\tilde{T}=t\mid X^{\prime}_{(1:t)\setminus k})+\delta,$ (16)
@@
We would like to emphasize that this relaxed $(\epsilon,\delta)$-differential privacy guarantee in (16) is not the same as the classical definition of $(\epsilon,\delta)$-DP in literature [12]. In the classical definition, the entire database—including the differing entry—is fixed, and the probability is taken only over the internal randomness of the algorithm, that is, $\mathbf{P}_{\mathcal{A}}(\mathcal{A}(D)\in\mathcal{S})\leq e^{\epsilon}\mathbf{P}_{\mathcal{A}}(\mathcal{A}(D^{\prime})\in\mathcal{S})+\delta$ for all neighboring datasets $D,D^{\prime}$ and measurable sets $\mathcal{S}$. In contrast, our slightly modified formulation (16) treats the differing entry itself as random and introduces a small probability $\delta$ to capture rare failure events where the LLR at differing entries exceeds a bounded threshold, i.e., $|\ell(X_{k})-\ell(X_{k}^{\prime})|>A_{\delta}$. Conditional on the high-probability event $|\ell(X_{k})|\leq A_{\delta}/2,|\ell(X_{k}^{\prime})|\leq A_{\delta}/2$, our procedure still achieves $\epsilon$-DP, and $\delta$ thus accounts for the probability of violating this condition. The same adaptation has also been used in [9] for handling unbounded LLRs in private change detection. To avoid confusion, we named our proposed procedure as $\delta$-DP-CUSUM procedure instead of $(\epsilon,\delta)$-DP-CUSUM procedure.
@@
For the general case with unbounded LLR, for any $\epsilon>0$, $\delta\in(0,1),$ $b>2,$ and the corresponding $A_{\delta}$ defined in (15), we have
@@
Then, for any $\epsilon>0$, $\delta\in(0,1),b>0,$ and the corresponding $A_{\delta}$ defined in (15), we have the worst-case average detection delay satisfies
@@
$b=b_{\gamma}=\frac{\log\gamma}{h(\epsilon,A_{\delta})}(1+o(1)),$ (19)
to guarantee the ARL constraint $\mathbf{E}_{\infty}[\tilde{T}]\geq\gamma$ is met. Under this choice of threshold, the resulting WADD of the $\delta$-DP-CUSUM procedure is upper-bounded by
@@
Our proposed procedure will achieve the first-order asymptotical optimality as the classical CUSUM procedure if $\epsilon\geq 2A_{\delta}.$ On the other hand, if $\epsilon<2A_{\delta}$, the upper bound becomes $\frac{\log\gamma}{I_{0}}(\frac{2A_{\delta}}{\epsilon})(1+o(1))$, which is greater than the optimal detection delay $\frac{\log\gamma}{I_{0}}(1+o(1))$ of classical CUSUM. Therefore, a smaller value of $h$ implies greater degradation in detection performance due to the privacy requirement. We further provide a Gaussian example below to illustrate how the privacy parameters $\epsilon$ and $\delta$ affect the detection efficiency of our proposed procedure by visualizing the value of $h(\epsilon,A_{\delta}).$ We emphasize that $\delta$ is user-specified and it yields the trade-off between privacy stringency and detection delay. In practice, one can choose the largest $\delta$ acceptable under one’s privacy requirements to make $h$ closer to 1, and thus reduce the detection delay.
@@
We provide an example under Gaussian distributions to demonstrate the theoretical effect of privacy parameters $\epsilon,\delta$ on detection delay. Assume one-dimensional Gaussian mean shift from the pre-change distribution $N(0,1)$ to the post-change distribution $N(\mu,1)$ with $\mu>0$. Therefore $f_{0}(x)=\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}x^{2}}$ is the pdf of pre-change distribution, and $f_{1}(x)=\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}(x-\mu)^{2}}$ is the pdf of post-change distribution. The LLR equals $\ell(X)=\log f_{1}(X)/f_{0}(X)=\mu X-\frac{1}{2}\mu^{2}$. Taking $A_{\delta}=2|\mu|\cdot z_{\delta/4}+\mu^{2}$ ensures $\max_{i=0,1}\,\mathbf{P}_{X\sim f_{i}}\left(2|\ell(X)|\geq A_{\delta}\right)\leq\delta/2$. The quantity $h(\epsilon,A_{\delta})=\min\{\frac{\epsilon}{2A_{\delta}},1\}$ is the effective privacy scaling factor that directly influences the asymptotic performance of the $\delta$-DP-CUSUM procedure. In particular, the best-case scenario corresponds to $h(\epsilon,A_{\delta})=1$, under which the $\delta$-DP-CUSUM procedure is guaranteed to be first-order asymptotically optimal. In contrast, smaller values of $h$ indicate a greater degradation in detection efficiency. Fig. 1 visualizes $h(\epsilon,A_{\delta})$ over a wide range of $\epsilon$ and $\delta$ values for different magnitude of the mean shift $\mu\in\{0.1,0.25,0.5\}$. The plots show that $h(\epsilon,A_{\delta})$ generally decreases as $\mu$ increases, because $A_{\delta}$ increases with $\mu$, thereby leading to a stronger impact of privacy on detection performance. The red dashed line in Fig. 1 serves as the boundary for optimal detection given by Theorem 2: any privacy parameters satisfying $\epsilon\geq 2A_{\delta}$ (equivalently $h(\epsilon,A_{\delta})=1$) guarantee that the $\delta$-DP-CUSUM procedure achieves first‑order asymptotic optimality.
@@
Remark 2.
From Fig. 1, we observe that stronger signals require larger values of $\epsilon$ and $\delta$ to attain optimality. Under the Gaussian example, this is because $A_{\delta}=2|\mu|\cdot z_{\delta/4}+\mu^{2}$ increases with $|\mu|$ and thus larger $\mu$ requires larger $\epsilon$ value to ensure optimality; conversely, under fixed $\mu$, smaller $\delta$ also leads to larger $A_{\delta}$ and thus requires larger $\epsilon$. It is worthwhile mentioning that this phenomenon is universal across general distributions. Intuitively, for stronger change signals (such as the case with larger $\mu$ in the Gaussian example), the LLR tends to take larger values with higher probability (and can become unbounded as assumed). Such a larger magnitude of LLR makes the CUSUM statistic more revealing of the raw data, necessitating stronger noise injection to preserve privacy—thereby degrading detection performance and requiring more relaxed privacy parameters $(\epsilon,\delta)$ for optimality. In contrast, for weaker change signals (such as smaller $\mu$ in the Gaussian setting), the LLR remains smaller and bounded with higher probability, requiring less noise and achieving asymptotic optimality even under smaller $(\epsilon,\delta)$. These observations also demonstrate the fundamental trade-off between detection delay and privacy gain: larger changes are easier to detect but require more effort to satisfy the DP guarantee.
@@ V-A DP-CUSUM with Bounded LLR: Laplace Distribution
We consider a mean shift in Laplace distribution family $\text{Lap}(\mu,1),$ which has a pdf $f(x|\mu)=\exp\left(-|x-\mu|\right)/2.$ Note we use a different notation here because we have used the shorthand notation $\text{Lap}(b)$ before for denoting the zero‑mean Laplace noise added in the detection procedure. We let the pre-change distribution be $\text{Lap}(0,1)$ and the post-change distribution be $\text{Lap}(\mu,1)$, with a nonzero post-change mean $\mu\neq 0$. Note the LLR between $\text{Lap}(\mu,1)$ and $\text{Lap}(0,1)$ is $\ell(x)=\left(|x|-|x-\mu|\right),$ which implies the sensitivity of the LLR is $\Delta=\sup_{x,x^{\prime}\in\mathbb{R}}|\ell(x)-\ell(x^{\prime})|=2|\mu|$.
@@
In Fig. 2(a), it can be seen that the detection delays for $\epsilon=0.8$ and $1$ are close to that of the exact CUSUM. This is consistent with our theoretical results on WADD in Theorem 3. Specifically, for $\mu=0.2$, we have $\Delta=2\mu=0.4$. By Eq. (14), the DP-CUSUM procedure is guaranteed to be asymptotically optimal when $\epsilon\geq 2\Delta=0.8$, which aligns with Fig. 2(a). As $\epsilon$ decreases, the delay increases, especially at larger ARL values, reflecting the trade-off between privacy and detection. A similar pattern is observed in Fig. 2(b) for $\mu=0.5$, where $\Delta=1$. It can be seen that the delay for $\epsilon=2\Delta=2$ matches the exact CUSUM, while smaller $\epsilon$ leads to worse performance.
@@
(a) $\text{Lap}(0,1)\to\text{Lap}(0.2,1)$ (b) $\text{Lap}(0,1)\to\text{Lap}(0.5,1)$
@@
We further compare the performance of our DP-CUSUM procedure with the baseline method, online private change-point detector (OnlinePCPD) [9], as shown in Fig. 3. OnlinePCPD builds on its offline Report Noisy Max framework and uses a fixed-size sliding window of width $w$ over the data stream. At each time step $t$, the algorithm considers a sliding window of recent samples $X_{t-w+1},\ldots,X_{t}$} and, for every candidate change-point $k\in\{t-w+1,\ldots,t\}$, computes the partial log-likelihood ratio $S_{k,t}=\sum_{i=k}^{t}\ell(X_{i})$. The statistic used for detection is $\tilde{S}_{t}=(\max_{t-w+1\leq k\leq t}S_{k,t})+Z_{t}$ with independent Laplace noise $Z_{t}\sim{\rm Lap}(8\Delta/\epsilon)$. The stopping rule is $\tilde{T}:=\inf\{t:\tilde{S}_{t}\geq b+W\}$, where $W\sim{\rm Lap}(4\Delta/\epsilon)$. This procedure is non-recursive, since all $S_{k,t}$ must be recomputed within each sliding window, resulting in an $O(w)$ per-step computational cost. While our proposed DP-CUSUM has only $O(1)$ per-step computational cost due to its recursive update of the detection statistics. Notably, in OnlinePCPD, Laplace noise $\text{Lap}(8\Delta/\epsilon)$ is added to the detection statistic and $\text{Lap}(4\Delta/\epsilon)$ to the threshold. In contrast, our method only adds $\text{Lap}(2\Delta/\epsilon)$ to both the statistic and the threshold, while still achieving the same level of $\epsilon$-DP, as guaranteed by Theorem 1. We adopt the same window size of 700 as used in [9]. Fig. 3 shows that the detection delay of OnlinePCPD is substantially larger than that of our DP-CUSUM procedure across all privacy regimes (even when $\epsilon\leq 2\Delta$), particularly when the ARL is large, under both mean shift scenarios $\mu=0.2$ and $\mu=0.5$ in the Laplace distribution setting.
@@
(a) $\text{Lap}(0,1)$ $\to$ $\text{Lap}(0.2,1)$ (b) $\text{Lap}(0,1)$ $\to$ $\text{Lap}(0.5,1)$
@@ V-B $\delta$-DP-CUSUM with Unbounded LLR: Normal Distribution
For $\mu=0.1$, we have $A_{\delta}=0.402$. By Eq. (20), the $\delta$-DP-CUSUM procedure is guaranteed to be asymptotically optimal when $\epsilon\geq 2A_{\delta}=0.804$, which aligns with the results in Fig. 4(a) where the detection delays for $\epsilon=1$ and $1.5$ are close to that of the exact CUSUM. A similar pattern is observed for $\mu=0.5$, where $A_{\delta}=2.21$. In Fig. 4(b), the delay for $\epsilon=4$ is close to that of the exact CUSUM, while a smaller $\epsilon$ leads to worse performance. It is also worth noting that performance degradation is less pronounced for smaller $\mu$ under the same $\epsilon$, likely because the detection problem is already difficult for smaller $\mu$ values—leading to larger delays even for the exact CUSUM procedure.
@@ VI Conclusion
We introduce a theoretically grounded DP-CUSUM procedure for sequential change detection with differential privacy constraints. We derive the lower bound to the average run length and the upper bound to the worst-case average detection delay of DP-CUSUM. The theoretical results also yield the first-order asymptotic optimality of DP-CUSUM under weak privacy requirements. Experimental results on synthetic data demonstrate the good performance of our DP-CUSUM procedure compared with the baseline method and demonstrate our theoretical findings. Our theoretical analysis provides analytical techniques for future extensions and the design of DP-CUSUM variants across diverse practical applications. Future work includes adapting this framework to more complex settings, such as unknown post-change distributions, multi-sensor detection with controlled sensing, network-structured data, and related applications.
Acknowledgments
The authors would like to thank the two anonymous reviewers and the Associate Editor for their insightful comments and constructive suggestions, which have greatly improved the clarity and quality of this work.
@@ Appendix A Useful Lemmas
$|S_{t}-S^{\prime}_{t}|$
$=|\max\left(0,S_{t-1}\right)+\ell(X_{t})-\max\left(0,S^{\prime}_{t-1}\right)-\ell(X_{t})|$
$\leq|S_{t-1}-S^{\prime}_{t-1}|\leq\Delta_{t-1}.$
@@
Lemma 4.
[18, 24] For the CUSUM statistics $S_{t}$ in (3), under the pre-change measure we have
@@
The proof can be found in [18, Lemma 3] and we omit here. ∎
@@
$\mathbf{E}_{\infty}[e^{\lambda S_{t}}]$
$=\int_{0}^{\infty}\mathbf{P}_{\infty}(e^{\lambda S_{t}}\geq y)dy=\int_{0}^{\infty}\mathbf{P}_{\infty}(S_{t}\geq\frac{\log y}{\lambda})dy$
$\overset{(i)}{=}\int_{-\infty}^{\infty}\mathbf{P}_{\infty}(S_{t}\geq u)\lambda e^{\lambda u}du$
$=\int_{-\infty}^{0}\mathbf{P}_{\infty}(S_{t}\geq u)\lambda e^{\lambda u}du+\int_{0}^{\infty}\mathbf{P}_{\infty}(S_{t}\geq u)\lambda e^{\lambda u}du$
@@
where $(i)$ by change of variables $y=e^{\lambda u}$, and $(ii)$ by $\mathbf{P}_{\infty}(S_{t}\geq u)\leq 1$ and $\mathbf{P}_{\infty}(S_{t}\geq u)\leq e^{-u}$.